OpenStudy (anonymous):

Solve the following systems of equations using the addition (elimination) method. What type of system is it? Name the solution if there is one. 5x=20+y 16=3y+4x

OpenStudy (anonymous):

5x -y =20 , 4x+3y = 16; multiply with 3 on both sides for equation 1 and add them , 15x-3y = 60 , 4x+3y= 16 => 19x=76 => x = 4 and so y = 0

OpenStudy (anonymous):

which is also inconsistent, I again suggest you to read the article and you should easily be able to solve such simpler equations